# BiLO: Bilevel Local Operator Learning for PDE Inverse Problems. Part II: Efficient Uncertainty Quantification with Low-Rank Adaptation

### v1

@@ 2.3 Review of Markov Chain Monte Carlo (MCMC) Methods
3:     Sample $\theta^{\prime}\sim Q(\cdot|\theta^{(k-1)})$
4:     Compute acceptance probability
$\alpha=\min\left(1,\exp\left[U(\theta^{(k-1)})-U(\theta^{\prime})\right]\cdot\frac{Q(\theta^{(k-1)}|\theta^{\prime})}{Q(\theta^{\prime}|\theta^{(k-1)})}\right)$
5:     Sample $u\sim U[0,1]$
6:     if $u<\alpha$ then
7:         Accept the proposal: $\theta^{(k)}\leftarrow\theta^{\prime}$
8:     else
9:         Reject the proposal: $\theta^{(k)}\leftarrow\theta^{(k-1)}$
10:     end if
11:end for
Hamiltonian Monte Carlo (HMC)
In HMC, an auxiliary momentum variable $\rho$ is introduced, and the Hamiltonian is defined as
$H(\theta,\rho)=U(\theta)+K(\rho)=-\log P(\theta|D)+\frac{1}{2}\rho^{T}M^{-1}\rho,$ (9)
where $M$ is the mass matrix, which is symmetric and positive definite. $M$ can be user-defined or adaptively learned in a warmup phase. HMC samples from the joint distribution $P(\theta,\rho)\propto\exp(-H(\theta,\rho))$ from the Hamiltonian dynamics.
$\frac{d\theta}{dt}$ $=M^{-1}\rho$ (10)
$\frac{d\rho}{dt}$ $=-\nabla_{\theta}U(\theta).$
There are many variants of HMC, and here we focus on the most popular choice, which involves the leapfrog integrator with a Metropolis-Hastings step. This can be viewed as an instance of the MH algorithm, where the proposal distribution is defined by the Hamiltonian dynamics. The scheme requires computing the gradient of the potential energy. Using the Hamiltonian dynamics, we can generate distant proposals with high acceptance rates. In sum, HMC is a powerful and successful sampling method [16].
Leapfrog Integrator
The Hamiltonian dynamics can be solved using numerical integrators. The most commonly used integrator is the leapfrog method, which is a symplectic integrator that preserves the Hamiltonian structure [16]. We sample a momentum variable $\rho$ from a Gaussian distribution $\mathcal{N}(0,M)$, and then simulate the Hamiltonian dynamics for $L$ steps with a fixed time step size $\delta t$. For $i=0,1,\dots,L-1$:
$\rho_{i+\frac{1}{2}}$ $=\rho_{i}-\frac{\delta t}{2}\nabla_{\theta}U(\theta_{i})$ (11)
$\theta_{i+1}$ $=\theta_{i}+\delta tM^{-1}\rho_{i+\frac{1}{2}}$
$\rho_{i+1}$ $=\rho_{i+\frac{1}{2}}-\frac{\delta t}{2}\nabla_{\theta}U(\theta_{i+1}),$
Both $\delta t$ and $L$ are hyperparameters that are predefined or adaptively tuned during a warm-up phase.
Metropolis-Hastings Step
Solving the Hamiltonian dynamics numerically introduces error, which can be corrected using the Metropolis-Hastings step [15]. Starting from some state $(\theta_{0},\rho_{0})$, we perform $L$ leapfrog steps to arrive at $(\theta_{L},\rho_{L})$. We then perform a Metropolis-Hastings step: we accept the proposal $\theta_{L}$ with probability $\alpha$, where $\alpha=\min\left(1,\exp\left(H(\theta_{0},\rho_{0})-H(\theta_{L},\rho_{L})\right)\right)$ If the proposal is accepted, we set $\theta^{(k)}=\theta_{L}$; otherwise, we set $\theta^{(k)}=\theta^{(k-1)}$. The momentum variable $\rho$ is discarded, and is resampled at the start of the next leapfrog step.
HMC Algorithm
We summarize the HMC algorithm in Algorithm 2. $\theta^{(k)}$ denotes the $k$-th sample, while $\theta^{(k)}_{i}$ denotes the PDE parameters at the $i$-th step of the leapfrog integrator for the $k$-th proposal.
Algorithm 2 Hamiltonian Monte-Carlo with Leapfrog Integrator
1:Initial states for $\theta^{(0)}$ and time step size $\delta t$
2:for $k=0,1,2,\dots,N$ do
@@
4:     $(\theta_{0},r_{0})\leftarrow(\theta^{(k)},r)$
5:     for $i=0,1,\dots,L-1$ do
6:         $r_{i+1/2}\leftarrow r_{i}-\frac{\delta t}{2}\nabla_{\theta}U(\theta_{i})$
7:         $\theta_{i+1}\leftarrow\theta_{i}+\delta tM^{-1}r_{i+1/2}$
8:         $r_{i+1}\leftarrow r_{i+1/2}-\frac{\delta t}{2}\nabla_{\theta}U(\theta_{i+1})$
9:     end for
10:     $\alpha\leftarrow\min\left(1,\exp\left(H(\theta_{0},r_{0})-H(\theta_{L},r_{L})\right)\right)$
11:     With probability $\alpha$, set $\theta^{(k)}\leftarrow\theta^{(k)}_{L}$; otherwise, set $\theta^{(k)}\leftarrow\theta^{(k-1)}$
12:end for
We note that sampling remains an active area of research. For example, many hyperparameters in HMC can be made adaptive. Further, there have been many advancements over HMC, such as the No-U-Turn Sampler (NUTS) [43], higher order integrator [44], or stochastic HMC [45] and we defer the use of such methods for future work. For simplicity and for comparison with [24], we take $M$ as the identity matrix, and we use the leapfrog method with fixed step size $\delta t$ and length $L$, which is also one of the most common versions of HMC. In the next section (2.4), we describe how to use HMC in the BiLO framework.
2.4 Algorithm and Theory of BiLO-HMC
Algorithm
The key component of our method is to compute the gradient of the potential energy $U(\theta)$. In BiLO, the potential energy depends on the weights of the neural network and is denoted as $U(\theta,W)$. At the optimal weights $W^{*}(\theta)$, the local operator $u(\mathbf{x},\theta;W^{*}(\theta))$, the potential energy $U(\theta,W*^{(}\theta))$, and its gradient $\nabla_{\theta}U(\theta,W*^{(}\theta))$ are expected to approximate the exact local operator $u(\mathbf{x},\theta)$, the true potential energy $U(\theta)$, and the true gradient $\nabla_{\theta}U(\theta)$, respectively. Therefore, each time $\theta$ is updated, we need solve the lower level problem.
In practice, we solve the lower level problem to some tolerance $\epsilon$ using our subroutine LowerLevelIteration in Algorithm 4. It is written as a simple gradient descent method, but in practice, we can use more advanced optimization methods such as Adam [46]. In the HMC, we call the subroutine LowerLevelIteration each time before computing the gradient of the potential energy $\nabla_{\theta}U(\theta,W)$. The complete BiLO-HMC algorithm is detailed in Algorithm 3.
Algorithm 3 BiLO-HMC
2:for $k=0,1,2,\dots,N$ do
3:     Sample $r\sim\mathcal{N}(0,M)$
4:     $(\theta_{0},r_{0})\leftarrow(\theta^{(k)},r)$
@@ 2.4 Algorithm and Theory of BiLO-HMC
15:     with probability $\alpha$, set $\theta^{(k)}\leftarrow\theta_{L}$; otherwise, set $\theta^{(k)}\leftarrow\theta^{(k-1)}$
@@
Theoretical Results
The accuracy of the leapfrog integrator depends on the accuracy of the potential energy gradient, $\nabla_{\theta}U(\theta,W)$, and the accuracy of the potential posterior distribution depends on the accuracy of the PDE solution. Both errors are attributed to the inexact solution of the lower-level optimization problem. The ideal optimal weights $W^{*}(\theta)$ satisfy the PDE for all $\theta$: $\mathcal{F}(u(\theta,W^{*}(\theta)),\theta)=0$. It’s practical approximation is denoted by $\overline{W}$, which is obtained by terminating the optimization once the lower-level loss is within a specified tolerance, $\mathcal{L}_{\text{LO}}(\theta,\overline{W})\leq\epsilon$. The true gradient of the upper-level objectives is $g_{true}(\theta)=d_{\theta}U(\theta,W^{*}(\theta))$. In BiLO, we compute an approximate gradient $g_{a}(\theta)=\nabla_{\theta}U(\theta,\overline{W})$, which does not consider the dependency of $\overline{W}$ on $\theta$. Under mild assumptions of PDE operator stability, smoothness, and Lipschitz continuity of the solution map, we demonstrated in Part I of our study [10] that $\norm{g_{a}-g_{true}}=O(\epsilon)$.
In addition to the error of the approximate gradient, inexact lower-level optimization also introduces a static error in the posterior distribution targeted by the BiLO-HMC sampler. Specifically, denote $u^{*}=u(\cdot,\theta;W^{*}(\theta))$, and $\bar{u}=u(\cdot,\theta;\overline{W})$. The practical posterior $\bar{\pi}(\theta)$ is based on inexact solution of the PDE $\bar{u}$, while the ideal posterior $\pi_{*}(\theta)$ is based on the exact solution $u^{*}$. In Theorem A.1, given in A, we quantify this static discrepancy by bounding the Kullback–Leibler divergence:
@@ 2.5.1 Pre-training and Fine-tuning
The additional data loss is not mandatory for training the local operator with fixed $\theta_{0}$. However, the additional data loss can speed up the training process [47], and is computationally inexpensive as we only need one numerical solution corresponding to $\theta_{0}$.
@@
One might want to keep a copy of $W(\theta^{(k-1)})$ and “backtrack” in case the proposal $\theta^{(k)}_{L}$ is rejected. We report that in practice, due to the high acceptance rate of HMC, this is not a problem. Even when $W(\theta^{(k)}_{L})$ is rejected, continuing the fine-tuning to recover $W(\theta^{(k-1)})$ is fast. We report that backtracking increases the memory use and does not noticeably improve the performance of the algorithm.
In Algorithm 4, we fine-tune all the weights of the neural network $W$. We call this approach full fine-tuning (Full FT). For large neural networks, this can still be computationally expensive. In the next section, we describe how to use the Low-Rank Adaptation (LoRA) technique [38] to reduce the number of trainable parameters and speed up the fine-tuning process.
2.5.2 Fine-tuning with Low Rank Adaptation (LoRA)
Low-Rank Adaptation (LoRA) is a technique initially developed for large language models [38]. In LoRA, instead of updating the full weight matrix, we learn an low-rank modification to the weight matrix. Below we describe the details of LoRA in the context of BiLO, following the notations used in Sec. 2.2.
@@ 2.5.2 Fine-tuning with Low Rank Adaptation (LoRA)
where $\theta$ are the PDE parameters computed in leapfrog. The full FT subroutine (Algorithm 4) can be replaced by the LoRA fine-tuning in Algorithm 5. In this work, we use the original LoRA [38] as a proof-of-concept and we note that more recent developments of LoRA might further increase efficiency and performance [48, 49, 50].
@@
In addition, with LoRA, the optimization problem is solved in a lower dimensional subspace of the full weight space. Therefore, it is expected that more iterations are needed to reach the same level of loss as full-model fine-tuning [51]. On the other hand, optimizing over a subspace can serve as a regularization [52]. In numerical experiments, we show that as the model size increases, the benefit outweighs the cost, and LoRA can significantly speed up the fine-tuning process.
2.6 Difference from BPINN
We briefly review the Bayesian PINN (BPINN) framework [24] and highlight the differences with BiLO.
2.6.1 Review of BPINN
Within the BPINN framework, the solution of the PDE is represented by a Bayesian neural network $u_{\rm PINN}(\mathbf{x},W)$, where $W$ denotes the weights of the neural network that are taken to be random variables. Note that $\theta$ is not an input to the neural network. For the Bayesian Neural Network, the prior distribution of the weights, $P(W)$, is usually assumed to be an i.i.d. Normal distribution.
The likelihood of the observation data $D_{u}=\{(\mathbf{x}_{u}^{(i)},\hat{u}^{(i)})\}_{i=1}^{N_{u}}$, where $\hat{u}^{(i)}=u(\mathbf{x}_{u}^{(i)})+\eta_{d}$, $\eta_{d}\sim\mathcal{N}(0,\sigma_{d}^{2})$, is given by
$P(D_{u}|W)=\prod_{i=1}^{N_{u}}\frac{1}{\sqrt{2\pi\sigma_{d}^{2}}}\exp\left(-\frac{1}{2\sigma_{d}^{2}}\left|u_{\rm PINN}(\mathbf{x}_{u}^{(i)},W)-\hat{u}^{(i)}\right|^{2}\right)$ (14)
Notice that the likelihood of the observation data $D_{u}$ only depends on the neural network weights $W$, and does not depends on the PDE parameters $\theta$ as BiLO does.
In BPINN, we denote the PDE as $\mathcal{G}(u(\mathbf{x}),\theta)=f(\mathbf{x})$, where $\mathcal{G}$ is the differential operator and $f$ is the forcing term. Given noisy measurements of the forcing term $D_{f}=\{(\mathbf{x}_{f}^{(i)},\hat{f}^{(i)})\}_{i=1}^{N_{f}}$, where $\hat{f}^{(i)}=f(\mathbf{x}_{f}^{(i)})+\eta_{f}$ and $\eta_{f}\sim\mathcal{N}(0,\sigma_{f}^{2})$, then
$P(D_{f}|W,\theta)=\prod_{i=1}^{N_{f}}\frac{1}{\sqrt{2\pi\sigma_{f}^{2}}}\exp\left(-\frac{1}{2\sigma_{f}^{2}}\left|\mathcal{G}(u_{\rm PINN}(\mathbf{x}_{f}^{(i)};W),\theta)-\hat{f}^{(i)}\right|^{2}\right)$ (15)
Assuming $\theta$ and $W$ are independent, the following joint posterior is sampled using HMC.
2.6.2 Challenges for BPINNs
One of the main differences between BPINN and BILO lies in the treatment of the neural network weights, $W$. For PDE inverse problems, the PDE parameters $\theta$ are usually low-dimensional. In the BPINN framework, the weights $W$ are treated as random variables that must be sampled from the joint posterior distribution alongside the PDE parameters $\theta$. However, $W$ is high-dimensional, making sampling challenging.
In contrast, the BiLO framework treats the weights $W$ as deterministic variables. For any given parameter $\theta$ in the sampling process, the optimal weights $W^{*}(\theta)$ are found through a direct optimization of the lower-level problem, as defined in Eq. (6). This approach avoids placing a prior on $W$ and more importantly bypasses the challenge of sampling from the high-dimensional and often complex posterior distribution of the network weights.
Modeling uncertainty in the forcing term $f$ can also be nuanced. In a BPINN, the solution $u$ is represented by a Bayesian neural network, and uncertainty originates from the prior distribution placed on the network weights, $P(W)$. Uncertainty over the weights $W$ then propagates through the differential operator $\mathcal{G}$ to induce a distribution on the model’s estimate of the forcing term, $f=\mathcal{G}(u(\cdot;W),\theta)$. This “top-down” uncertainty propagation from the model’s prior contrasts with cases where it is more physically meaningful for uncertainty in the forcing term $f$ itself (e.g., from noisy measurements or an explicit prior, $P(f)$) to propagate to the solution $u$.
Another nuance lies in the interpretation of $\sigma_{f}$ and $D_{f}$. When $D_{f}$ represents noisy physical measurements of the forcing term $f$, and $\sigma_{f}$ is the noise level, the likelihood $P(D_{f}|W,\theta)$ allows one to incorporate the PDE. However, if no such measurements are available, then the physics-informed component of the likelihood vanishes. On the other hand, if there is no noise in $D_{f}$, then $\sigma_{f}$ is 0, and the likelihood becomes singular and cannot be sampled by most sampling methods. An alternative interpretation is to view $D_{f}$ and $\sigma_{f}$ as user-defined constructs. Similar to the residual loss in PINNs, they serve as a soft PDE constraint. In this case, $\sigma_{f}$ plays the role of a penalty parameter: smaller values lead to more accurate solution of the PDE.
Regardless of the interpretation, having a small $\sigma_{f}$ is computationally demanding. This challenge stems from the stability requirements of HMC, which dictates that the leapfrog step size, $\delta t$, is limited by the inverse square root of the potential energy’s maximum curvature [16]. In the BPINN framework, this maximum curvature is proportional to $\lambda_{\rm max}/\sigma_{f}^{2}$, where $\lambda_{\rm max}$ is the largest eigenvalue of the Hessian of the unscaled residual loss, $\mathcal{L}_{\rm res}$. This unscaled loss is often ill-conditioned, with $\lambda_{\rm max}$ values known to exceed $10^{3}$ [53]. Consequently, applying the HMC stability limit imposes a scaling law on the step size, forcing $\delta t$ to be of order $\mathcal{O}(\sigma_{f}/\sqrt{\lambda_{\rm max}})$. In contrast, the BiLO framework does not require sampling the weights $W$, and thus avoids the stability limit imposed by the Hessian of the residual loss. This allows BiLO to use larger step sizes, leading more distant proposals and more efficient sampling.
3 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance of the BiLO framework with LoRA for Bayesian inverse problems. We mainly consider the following three methods for comparison:
  • Reference: The PDE is solved with high accuracy using numerical methods on a fine grid, are thus considered as “exact” solution of the PDE. The posterior distribution of the PDE parameters is sampled using Metropolis-Hastings (MH) with 100,000 steps. For simplicity, we use the transitioin $\theta^{\prime}=\theta+\eta$, where $\eta$ is uniformly distributed in $[-\Delta\theta,\Delta\theta]$. $\Delta\theta$ is chosen so that the acceptance rate is around 0.3 [54]. Due to the accurate PDE solution, the obtained posterior distribution is considered as the reference.
  • BPINN ($\sigma_{f}$): The PDE solution is represented by BPINN. The certain $\sigma_{f}$, is specified. The joint distribution of the PDE parameters and the neural network weights are sampled with HMC. This is the implementation in [24].
  • BiLO: The PDE solution is represented by BiLO, and the posterior distribution is obtained through HMC. This uses our algorithm 3 with full FT (Algorithm 4) by default.
We don’t explicitly mention HMC in the legends of the figures as it is used in both BPINN and BiLO. When demonstrating the efficiency of LoRA, we consider the following two different fine-tuning methods:
  • Full FT: full fine-tuning using Algorithm 4.
  • LoRA ($r$): LoRA fine-tuning with rank $r$, as in Algorithm 5.
In Section 3.1, we compare BiLO with BPINN on a nonlinear Poisson problem to demonstrate the advantages of BiLO: physical uncertainty quantification, stable sampling with larger step sizes, and reduced computational cost. In Section 3.2, we apply BiLO to infer stochastic rates from particle data, showcasing its ability to handle PDEs with singular forcing and discontinuous derivatives. We demonstrate the efficiency of LoRA. In Section 3.3, we solve a 2D Darcy flow problem with a spatially varying diffusion coefficient. Additional results and details of the numerical experiments can be found in B.
3.1 Comparison with BPINNs
We consider the 1D nonlinear Poisson equation [24]. Let $u^{\rm GT}(x)=\sin^{3}(6x)$, $\lambda=0.01$ and $k^{\rm GT}=0.7$. The boundary value problem is given by
$\lambda u_{xx}(x)+k\tanh(u(x))=f(x),\quad x\in(-0.7,0.7)$ (17)
$u(x)=\sin^{3}(6x)~{}{\rm at}~{}x=\pm 0.7,$ (18)
where
$f(x)=-\lambda u^{\rm GT}_{xx}(x)-k^{\rm GT}\tanh(u^{\rm GT}(x)).$ (19)
The goal is to infer $k$ from observations of $u$. We assume the prior of $k$ is the uniform distribution on $[0.4,1]$.
The “exact” solutions for different $k$ are obtained by solving the nonlinear system with finite difference discretizations using Newton’s method with initial guess $u^{\rm GT}$. The solution to the PDE is odd: both $u$ and $-u(-x)$ solve the PDE, and thus $u(0)=0$ for all $k$ in the range. Therefore, there should be no uncertainty at $u(0)=0$.
We compare BiLO and BPINN under different values of $\sigma_{f}$ ($10^{-2}$, $10^{-1}$, and $1$) in Figure 2. We begin by examining the inference results for the solution $u$. Figures (a), (b), (d), (e), and (f) show the ground truth solution $u^{\rm GT}$ (dashed line), noisy data (dots), the posterior mean (solid line), and the 95% confidence interval (shaded region), computed as the posterior mean ±$\pm 2$ standard deviations (std) of the inferred $u$. These plots correspond to the reference method, BiLO, and BPINN with $\sigma_{f}=1$, $0.1$, and $0.01$, respectively. In the reference result, $u^{\rm GT}$ lies entirely within the 95% confidence interval, the posterior mean closely matches $u^{\rm GT}$, and the uncertainty vanishes at $u(0)=0$, as expected. BiLO yields results that closely match the reference.
Figures (d), (e), and (f) present results from BPINN-HMC with $\sigma_{f}=1$, $0.1$, and $0.01$. With $\sigma_{f}=0.01$, the BPINN solution is visually close to the reference. At $\sigma_{f}=0.1$, however, the posterior mean fails to pass through the origin, and there is nonzero uncertainty at $u(0)=0$, which is unphysical and indicates an inaccurate PDE solution. With $\sigma_{f}=1$, the results degrade further: the confidence interval is overly broad, and the posterior mean simply interpolates the noisy data.
Figure 2(c) shows the probability density functions (PDFs) of the inferred parameter $k$ obtained via kernel density estimation (KDE). The PDFs from BiLO and BPINN with $\sigma_{f}=0.01$ are both close to the reference. In contrast, BPINN with $\sigma_{f}=0.1$ and $1$ yields inaccurate posteriors for $k$, with the PDF being overly broad and shifted away from the reference.
Overall, BiLO provides accurate results without needing to tune hyperparameters such as $\sigma_{f}$. In BPINN, larger values of $\sigma_{f}$ weaken the PDE constraint, leading to inaccurate posteriors for both the PDE parameters and the solution $u$. While BPINN can achieve accurate inference with sufficiently small $\sigma_{f}$, we show in the next figure that this comes at a significantly higher computational cost.
Figure 2: Inference results for the nonlinear Poisson problem. Noisy data, GT solution, mean and std of inferred $u$ using (a) Reference, (b) BiLO, (d-f) BPINN with varying $\sigma_{f}$. Panel (c) shows the probability density function (PDF) of inferred $k$. The soft PDE constraint of BPINN leads to unphysical uncertainty at $u(0)=0$ and larger uncertainty in $k$.
In Fig. 3, we compare the performance of BiLO and BPINN with $\sigma_{f}$ = 0.01 and 0.1 ($\sigma_{f}=1$ is not plotted due to its low accuracy for BPINN) under varying leapfrog step sizes $\delta t$ = $10^{-4}$, $10^{-3}$ and $10^{-2}$, with a fixed trajectory length $L=10$ for a fixed $N=5000$ (number of proposals). The effective sample size (ESS) [55] measures the number of independent samples in a Markov chain. Fig. 3(a) shows the acceptance rate, while Fig. 3(b) reports the ESS per wall time (in seconds), which reflects the overall efficiency of the sampler. In general, for fixed $L$, increasing $\delta t$ leads to longer trajectories and more decorrelated proposals, but also increases the risk of instability and rejection. As discussed in Sec. 2.6 and illustrated in Fig. 2, for BPINN, small $\sigma_{f}$ is required for accurate solution, but incurs a large computational cost as the step size $\delta t$ must be small to maintain stability. This is illustrated in Fig. 3 (a), where the acceptance rate of BPINN for fixed $\sigma_{f}$ drops significantly as $\delta t$ increases. Alternatively, for example, for fixed $\delta t=0.01$, reducing $\sigma_{f}$ from $0.1$ to $0.01$ leads to a dramatic drop in acceptance rate. For fixed $\delta t=0.1$, $\sigma_{f}=0.1$ has a near zero acceptance rate and $\sigma_{f}=0.01$ has zero acceptance rate.
In contrast, BiLO remains stable across a broad range of $\delta t$ and maintains high acceptance rates. This robustness arises from BiLO’s lower-dimensional sampling space, which avoids the high-dimensional weight space of neural networks and allows for efficient exploration of the space of the parameters with larger step sizes. By using larger $\delta t$ that is unattainable for BPINN with small $\sigma_{f}$, BiLO achieves higher ESS per wall time, as shown in Figure 3 (b).
Figure 3: Sampling efficiency of BiLO and BPINN for the nonlinear Poisson problem under varying leapfrog step sizes $\delta t\in\{10^{-4},10^{-3},10^{-2}\}$ with fixed trajectory length $L$. (a) Acceptance rate. (b) Effective sample size (ESS) per second. BPINN is sensitive to both $\delta t$ and $\sigma_{f}$, with smaller $\sigma_{f}$ (stronger PDE constraint) requiring smaller $\delta t$ for numerical stability, leading to reduced acceptance and efficiency. In contrast, BiLO remains robust across all tested $\delta t$, benefiting from its lower-dimensional parameter space and achieving significantly higher sampling efficiency.
3.2 Inferring Stochastic Rates from Particle Data
We consider a problem motivated by inferring the dynamics of gene expression from static images of mRNA molecules in cells [56, 57]. The steady-state spatial distribution of these molecules can be modeled by the following boundary value problem with a singular source:
Here, the solution $u(x)$ represents the intensity of a spatial Poisson point process describing the locations of mRNA particles in a simplified 1D domain. The parameter $\lambda$ is the dimensionless birth rate (transcription) of mRNA at a specific gene site $z$, while $\mu$ is a a degradation rate, and the boundary condition describes export across the nuclear boundary.
Given $M$ snapshots of mRNA locations $\{q_{i}^{j}\}$ from different cells, the goal is to infer the kinetic parameters $\lambda$ and $\mu$. Since the particle locations follow a Poisson point process, we can construct a likelihood function directly from their positions. The resulting log-likelihood is [56]:
[…]
We use this example to show the effect of LoRA. In Fig. 4(a), we show the 90% highest posterior density region (HDPR) and the maximum a posteriori (MAP) of the inferred parameters $\lambda$ and $\mu$ using the reference method and BiLO with LoRA rank 2, rank 4, and Full FT. In Fig. 4(b), we show the ground truth (GT), prior distributions, and posterior distributions of $\lambda$ and $\mu$ (the ESS are around 150). Table 2 reports the relative error of the MAP estimates and the area of the 90% HDPR with respect to the reference. All methods yield less than 4% relative error across all three metrics. Across fine-tuning strategies, both $\lambda_{\rm MAP}$ and $\mu_{\rm MAP}$ are slightly overestimated relative to the reference, which can be attributed to the inexact PDE solution. The LoRA rank 2 and rank 4 results are comparable to Full FT, despite requiring significantly fewer trainable parameters. Minor differences between methods are likely due to variability from finite MCMC sampling.
Figure 4: Inference of gene expression parameters $\lambda$ (birth rate) and $\mu$ (switch rate) using BiLO. (a): 90% highest density posterior region (HDPR) and maximum a posteriori (MAP) estimates obtained using BiLO with LoRA rank 2, rank 4, full FT, and reference. All results of BiLO are close to the reference solution. LoRA rank 2 and 4 yield comparable results to full FT. (b): Ground truth (GT), prior, and posterior distributions for $\lambda$ and $\mu$, with posterior sampled via BiLO using LoRA rank 2.
3.3 Darcy Flow Problem
We consider the following 2D Darcy flow problem on the unit square $\Omega=[0,1]\times[0,1]$:
$\begin{cases}-\nabla\cdot(D(\mathbf{x})\nabla u)=f(\mathbf{x})&\text{in }\Omega\\ u=0&\text{on }\partial\Omega\end{cases}$ (21)
where $f(\mathbf{x})=10$, $D(\mathbf{x})$ is unknown and needs to be inferred.
3.3.1 Inferring Unknown Functions

### v2

@@ 4.2 Algorithm
4:  $(\theta_{0},r_{0})\leftarrow(\theta^{(k-1)},r)$
@@
15:  with probability $\alpha$, set $\theta^{(k)}\leftarrow\theta_{L}$; otherwise, set $\theta^{(k)}\leftarrow\theta_{0}$
@@
4.3 Theoretical analysis
The accuracy of the leapfrog integrator depends on the accuracy of the potential energy gradient, $\nabla_{\theta}U(\theta,W)$, and the accuracy of the potential posterior distribution depends on the accuracy of the PDE solution. Both errors are attributed to the inexact solution of the lower-level optimization problem.
The ideal optimal weights $W^{*}(\theta)$ satisfy the PDE for all $\theta$: $\mathcal{F}[u(\cdot,\theta,W^{*}(\theta)),\theta]=\mathbf{0}$. Its practical approximation is denoted by $\overline{W}$, which is obtained by terminating the optimization once the lower-level loss is within a specified tolerance, $\mathcal{L}_{\text{LO}}(\theta,\overline{W})\leq\epsilon$. The true gradient of the upper-level objectives is $g_{true}(\theta)=d_{\theta}U(\theta,W^{*}(\theta))$. In BiLO, we compute an approximate gradient $g_{a}(\theta)=\nabla_{\theta}U(\theta,\overline{W})$, which does not consider the dependency of $\overline{W}$ on $\theta$. Under mild assumptions of PDE operator stability, smoothness, and Lipschitz continuity of the solution map, we proved that $\norm{g_{a}-g_{true}}=O(\epsilon)$ [zhangBiLOBilevelLocal2025a].
In addition to the error of the approximate gradient, inexact lower-level optimization also introduces a static error in the posterior distribution targeted by the BiLO-HMC sampler. Specifically, denote $u^{*}=u(\cdot,\theta;W^{*}(\theta))$, and $\bar{u}=u(\cdot,\theta;\overline{W})$. The practical posterior $\bar{\pi}(\theta)$ is based on inexact solution of the PDE $\bar{u}$, while the ideal posterior $\pi_{*}(\theta)$ is based on the exact solution $u^{*}$. In Theorem 1, given in the SM, we quantify this static discrepancy by bounding the Kullback–Leibler divergence:
@@ 4.4 Efficient Sampling with Low-Rank Adaptation (LoRA)
The additional data loss is not mandatory for training the local operator with fixed $\theta_{0}$. However, the additional data loss can speed up the training process [krishnapriyanCharacterizingPossibleFailure2021], and is computationally inexpensive as we only need one numerical solution corresponding to $\theta_{0}$.
@@
In Algorithm 2, we fine-tune all the weights of the neural network $W$. We call this approach full fine-tuning (Full FT). For large neural networks, this can still be computationally expensive. In the next section, we describe how to use the Low-Rank Adaptation (LoRA) technique [huLoRALowRankAdaptation2021] to reduce the number of trainable parameters and speed up the fine-tuning process.
Efficient Fine-tuning with Low Rank Adaptation (LoRA)
Low-Rank Adaptation (LoRA) is a technique initially developed for large language models [huLoRALowRankAdaptation2021]. In LoRA, instead of updating the full weight matrix, we learn an low-rank modification to the weight matrix. Below we describe the details of LoRA in the context of BiLO, following the notations used in Sec. 4.1.
@@
where $\theta$ are the PDE parameters computed in leapfrog. The full FT subroutine (Algorithm 2) can be replaced by the LoRA fine-tuning in Algorithm 3. In this work, we use the original LoRA [huLoRALowRankAdaptation2021] as a proof-of-concept and we note that more recent developments of LoRA might further increase efficiency and performance [zhangAdaLoRAAdaptiveBudget2023, wangLoRAGALowRankAdaptation2024, hayouLoRAEfficientLow2024].
@@
In addition, with LoRA, the optimization problem is solved in a lower dimensional subspace of the full weight space. Therefore, it is expected that more iterations are needed to reach the same level of loss as full-model fine-tuning [chenCELoRAComputationEfficientLoRA2025]. On the other hand, optimizing over a subspace can serve as a regularization [bidermanLoRALearnsLess2024]. In numerical experiments, we show that as the model size increases, the benefit outweighs the cost, and LoRA can significantly speed up the fine-tuning process.
4.5 Inferring Unknown Functions